OpenStudy (anonymous):

What is the standard deviation of the data set? 14 34 23 50 25 16 48 a. about 35.6 b. about 26.8 c. about 14.5 d. about 13.4 my answer: c. about 14.5 or is it d?? idk please help i got confused... on the scientific calculator, is Sx the symbol for standard deviation? or is it something else?

OpenStudy (cwtan):

Erm I duno what type of ur calculator are But My calculator show the ans is D

OpenStudy (cwtan):

Mayb we should do it manually to find it.

OpenStudy (anonymous):

my calculator is TI-84 plus texas instruments calculator.... what is the symbol or standard deviation on a calculator?

OpenStudy (kropot72):

@iheartfood Your calculation agrees with mine. Sx is correct for calculation of an estimate of the Standard Deviation based on a small sample. Lower case sigma is the symbol for the population standard deviation.

OpenStudy (anonymous):

So the answer is C 14.5 right??? @kropot72

OpenStudy (cwtan):

Standard Deviation= sqrt(total of x^2/n- mean^2) Mean =30 as i counted Total of x^2 =7566 7566/7=approx 1080 1080-900=180 sqrt 180=approx 13.41

OpenStudy (kropot72):

The answer is 14.5 (rounded to one decimal place because it is only an estimate of the population standard deviation).

OpenStudy (anonymous):

okay, thx

OpenStudy (kropot72):

You're welcome. Note: There is often confusion as to how to estimate the standard deviation of a small sample when details of the parent distribution are not known.

OpenStudy (kropot72):

On second thoughts, the question does not state that the data is from a sample. Therefore in this case the sample standard deviation calculation is not appropriate. The standard deviation is choice d. "about 13.4"

OpenStudy (anonymous):

the answer is D? is this according to the formula/equation @kropot72

OpenStudy (kropot72):

According to the calculated square root of the variance\[\sigma ^{2}=\frac{\Sigma[(X-Xbar)^{2}]}{n}\]

OpenStudy (anonymous):

oh okay thanks :)